OpenStudy (anonymous):

find the roots. 2x^2+4x-5=0 a.-2+- squrt 10/2 b.-2 +-sqrt 14/2 c.-4+- sqrt 14/4

OpenStudy (anonymous):

b

OpenStudy (anonymous):

your discriminant is b^2-4ac=4^2-4*2*(-5)=16+40=56 x1,2=(-b+/-sqrt(D))2a=(-4 +/- sqrt56)/4 = (-4 +/- 2sqrt14)/2 = (-2 +/- sqrt14)/2